Net Present Value Method and Present Value Index

Diamond and Turf Inc. is considering an investment in one of two machines. The sewing machine will increase productivity from sewing 110 baseballs per hour to sewing 198 per hour. The contribution margin per unit is $0.36 per baseball. Assume that any increased production of baseballs can be sold. The second machine is an automatic packing machine for the golf ball line. The packing machine will reduce packing labor cost. The labor cost saved is equivalent to $22 per hour. The sewing machine will cost $169,200, have an eight-year life, and will operate for 1,400 hours per year. The packing machine will cost $94,800, have an eight-year life, and will operate for 1,200 hours per year. Diamond and Turf seeks a minimum rate of return of 15% on its investments.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Determine the net present value for the two machines. Use the table of present values of an annuity of $1 above. Round to the nearest dollar.

Sewing Machine Packing Machine
Present value of annual net cash flows $ $
Amount to be invested $ $
Net present value $ $

b. Determine the present value index for the two machines. If required, round your answers to two decimal places.

Sewing Machine Packing Machine
Present value index
Answer #1
Annual cash flows:
Sewing Machine 44352 =(198-110)*0.36*1400
Packing Machine 26400 =22*1200
a
Sewing Machine Packing Machine
Present value of annual net cash flows 199007 118457
Amount to be invested 169200 94800
Net present value 29807 23657
b
Sewing Machine Packing Machine
Present value index 1.18 1.25
Workings:
Sewing Machine Packing Machine
Present value of annual net cash flows =44352*4.487 =26400*4.487
Sewing Machine Packing Machine
Present value index =199007/169200 =118457/94800